OpenStudy (anonymous):

What is the simplified form of the quantity 4 x squared minus 25 over the quantity 2 x plus 5 ? Answer 2x – 5, with the restriction x ≠ –five over 2 2x + 5, with the restriction x ≠ five over 2 2x – 5, with the restriction x ≠ five over 2 2x + 5, with the restriction x ≠ –five over 2

OpenStudy (anonymous):

ok so what bout the restriction

OpenStudy (phi):

The restriction exists because you are dividing by 2x-5 if x= 5/2 then 2x-5 becomes (put 5/2 in for x) 2*5/2-5= 5-5=0 You are not allowed to divide by 0, so do not allow x= 5/2 restrict x: x≠ 5/2